OpenStudy (anonymous):

Simplify 14m-21 __________ 7

OpenStudy (anonymous):



OpenStudy (anonymous):

reduce those fractions

OpenStudy (bibby):

If that doesn't help, factor the top half, like so: \(\huge \frac{14m-21}{7}=\frac{?(?-?)}{7}\)

OpenStudy (anonymous):

Yes, what bibby says is correct